OpenStudy (anonymous):

what values for theta(0≤theta≤2pi) satisfy the equation? 2sinthetacostheta+sqrt 3 theta

OpenStudy (freckles):

is there an equation?

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

ok so you meant to put =0 at the end

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I'm really confused x.x why r practice packets so hard >:C

OpenStudy (freckles):

and the other cos thingy but yeah anyways you can factor the left hand side \[2 \sin(\theta)\cos(\theta)+\sqrt{3} \cos(\theta)=0 \\ \text{ factor out the } \cos(\theta) \\ \cos(\theta)[2 \sin(\theta)+\sqrt{3}]=0 \]

OpenStudy (freckles):

set both factors =0 and solve for theta

OpenStudy (freckles):

\[\cos(\theta) \text{ or } \sin(\theta) =\frac{-\sqrt{3}}{2}\] use unit circle

OpenStudy (anonymous):

??!!

OpenStudy (freckles):

forgot the =0 part :p

OpenStudy (freckles):

\[\cos(\theta)=0 \text{ or } \sin(\theta) =\frac{-\sqrt{3}}{2}\] use unit circle *

OpenStudy (anonymous):

??? wits the answer suppose to be :( explanation too pls

OpenStudy (freckles):

sin(theta) is the y-coordinate on the unit circle find for what theta's you have the y value ( the second value) of the pair is -sqrt(3)/2

OpenStudy (freckles):

hint: look in quadrants 3 and 4

OpenStudy (anonymous):

okie :( wait the answer is B??

OpenStudy (freckles):

sin(3pi/4) is not -sqrt(3)/2 it is sqrt(2)/2

OpenStudy (anonymous):

so theres no answer? ._.

OpenStudy (freckles):

i didn't say that i just said the option you chose is incorrect because sin(3pi/4) is not -sqrt(3)/2

OpenStudy (freckles):

did you look at the unit circle you are looking for when the y-coordinate is -sqrt(3)/2

OpenStudy (anonymous):

D

OpenStudy (freckles):

the hint I gave is to look in the 3rd and 4th quadrant

OpenStudy (anonymous):

d??

OpenStudy (freckles):

https://www.mathsisfun.com/geometry/images/circle-unit-radians.gif look at this picture do you see the ordered pairs?

OpenStudy (freckles):

look at the second number of each pair you are looking to for which theta you have the y coordinate is -sqrt(3)/2

OpenStudy (freckles):

can you tell me looking at that chart thing where you see y coordinate is -sqrt(3)/2

OpenStudy (freckles):

ok look at that chart do you see under sin -sqrt(3)/2 ?

OpenStudy (freckles):

find it then use your finger and move it horizontally from that number to the theta number

OpenStudy (anonymous):

OH I GET WUT MEAN ITS A DUH ITS 4PI/3 IM SO STUPID

OpenStudy (freckles):

yes sin(4pi/3)=-sqrt(3)/2

OpenStudy (freckles):

and sin(5pi/3)=-sqrt(3)/2

OpenStudy (freckles):

and of course cos(theta)=0 is easy to solve and all of the choices included its answers which was pi/2 and 3pi/2

OpenStudy (anonymous):

:D thnx so much here a medal

OpenStudy (freckles):

np